OpenStudy (anonymous):

how do i simplify 27d/ 27d^2+18d

OpenStudy (ahsome):

Is this it: \[\frac{27d}{27d^2}+18d\] ?

OpenStudy (ahsome):

If it is, firstly, eliminate \(27d\) from the numerator and denominator \[\frac{27d}{27d^2}+18d\] \[(\frac{27d}{27d^2})/27d+18d\] \[\frac{1}{d}+18d\] \[\frac{1}{d}+\frac{18d}{1}\] Multiply the \(\frac{18d}{1}\) by \(d\) to get a common denominator \[\frac{1}{d}+(\frac{18d}{1})*\frac{d}{d}\] \[\frac{1}{d}+\frac{18d^2}{d}\] \[\frac{18d^2+1}{d}\]

OpenStudy (anonymous):

no it was what i had, but i was looking at it again and i think i got it. thanks any way. @Ahsome

OpenStudy (ahsome):

No problem @Manuela100 :)